# A discrete Consensus-Based Global Optimization Method with Noisy Objective Function

### v1

@@ 3.5 Convergence in expectation
$\mathbb{P}\left\{|{\cal I}_{\bar{k},r}|>0\right\}\geq 1-\delta\enspace\text{with}\ \delta\leq\hat{\delta}e^{-4\alpha M_{f}}$
where $\bar{k}$ is given in (56), $M_{f}$ is the upper bound from Corollary 3.5, and $\hat{\delta}\in(0,e^{4\alpha M_{f}})$. Then, for $\alpha$ sufficiently large and $M_{v}$ sufficiently small it holds ${\mathbb{E}}[V_{k_{*}}]\leq\varepsilon$, with $k_{*}$ given in (57). Moreover, ${\mathbb{E}}[V_{k}]\leq\mu^{k}{\mathbb{E}}[V_{0}]$ for $k=0,\dots,k_{*},$ with $\mu$ given in (55).
@@ 4 Numerical Results
In this section we investigate on the sensitivity of our approach to noise in the objective function evaluation. We first consider the minimization of the well known Rastrigin function with Gaussian noise on the objective function and then we analyze the behaviour of CBO applied to finite sum minimization problems employing subsampling estimators of the objective function rather than the exact objective function.
@@ 4.1 Rastrigin Function
We can observe that the absolute Gaussian noise affects the quality of the solution as expected, but we manage to obtain an average error of the order of $10^{-3}$ (against $10^{-5}$ obtained in the noise-free case), even when an high noise level is present. We also underline that the method is quite insensitive to this type of noise as increasing $\sigma_{0}$ we observe only a slight decrease of the overall obtained accuracy. Regarding the relative Gaussian noise, we note that the noise does not affect the accuracy of the computed approximation for values of the noise level smaller or equal than $0.25$, while the method is not able to handle higher values of the noise. In case of the mixed noise we obviously observe the effect of both the absolute and relative noise. Overall the observed behaviour is consistent with our theoretical analysis. Indeed, under Assumption 5, Theorems 3.15 and 3.16 shows that global convergence is obtained provided that the the upper bound $t_{0}+t_{1}M_{f}^{2}$ on the noise level is sufficiently small. In case of relative error, the bound on $\sigma_{1}$ is more severe than the bound on $\sigma_{0}$ at particles $x$ at which $f$ takes large value.
@@ 4.2 Subsampling
For both choices of the parameters we can observe that the final accuracy is not greatly affected by the subsampling, up to $\ell=0.25$ in the first case and $\ell=0.05$ in the second case. For $\gamma=0.1$ and $\xi=0.0056$ (Table 5) the number of iterations is approximately constant for all considered values of $\ell$, which implies that the computational cost decreases significantly as the sampling size decreases. For $\ell=0.5$ the overall computational cost is halved, but the average accuracy is less than $0.5\%$ smaller. For $\gamma=0.01$ and $\xi=0.1$ (Table 6) the number of iterations increases as the sampling factor decreases, but the overall computational cost decreases. The method that employs $\ell=0.1$ is approximately 2.5 times less expensive than the method that uses the full sample, while the average accuracy is only around $1\%$ smaller.
@@ Acknowledgements
The research that led to the present paper was partially supported by INDAM-GNCS through Progetti di Ricerca 2023 and by PNRR - Missione 4 Istruzione e Ricerca - Componente C2 Investimento 1.1, Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN) funded by the European Commission under the NextGeneration EU programme, project “Advanced optimization METhods for automated central veIn Sign detection in multiple sclerosis from magneTic resonAnce imaging (AMETISTA)”, code: P2022J9SNP, MUR D.D. financing decree n. 1379 of 1st September 2023 (CUP E53D23017980001), project “Numerical Optimization with Adaptive Accuracy and Applications to Machine Learning”, code: 2022N3ZNAX MUR D.D. financing decree n. 973 of 30th June 2023 (CUP B53D23012670006).
@@
[13, Lemma 3.3 ] If $M\in\mathbb{R}^{N\times N}$ is such that $\sum_{j=1}^{N}M_{ij}=m$ for any index $i=1,\dots,N$, then for every vector ${y}\in\mathbb{R}^{N}$ we have $\operatorname{diam}(My)\leq(m-\operatorname{erg}(M))\operatorname{diam}(y).$
@@
[13, Lemma 4.1] Given any matrix $M\in\mathbb{R}^{N\times N}$, the ergodicity coefficient satisfies the following inequality:
$\operatorname{erg}(M)\geq-2\|M\|_{\infty}$ for any $M\in\mathbb{R}^{N\times N}$
@@
$(M^{s}_{k})_{ij}=\begin{cases}1-\gamma-\eta^{i}_{k,s}+(\gamma+\eta^{i}_{k,s})\hat{v}^{\alpha}_{k,j}&\text{if }i=j\\ (\gamma+\eta^{i}_{k,s})\hat{v}^{\alpha}_{k,j}&\text{if }i\neq j\end{cases}$ (72)
@@
Replacing (74) and (75) into (73) we get the thesis. ∎
@@
$Q_{k,s}=\gamma\sum_{l=0}^{k}(x^{i}_{l,s}-\hat{x}^{\alpha}_{l,s}),\enspace P_{k,s}=\sum_{l=0}^{k}(x^{i}_{l,s}-\hat{x}^{\alpha}_{l,s})\eta_{l,s}^{i}$
@@
Since $S_{l,s}$ converges to $1-{\bf\vartheta}$ almost surely as $l$ goes to $\infty$, there exist two stochastic variables $C_{1},C_{2}>0$ such that
@@
$={\mathbb{E}}\left[\sum_{l=0}^{k}(x^{i}_{l,s}-\hat{x}^{\alpha}_{l,s})^{2}(\eta_{l,s}^{i})^{2}+\sum_{l=0}^{k-1}\sum_{\begin{subarray}{c}p=0\\ p\neq l\end{subarray}}^{k-1}(x^{i}_{l,s}-\hat{x}^{\alpha}_{l,s})(x^{i}_{p,s}-\hat{x}^{\alpha}_{p,s})\eta_{l,s}^{i}\eta_{p,s}^{i}\right]$
@@
Since $\lim_{l\rightarrow+\infty}S_{l,s}=1-{\bf\vartheta}>0$ almost surely, then there exists a constant $\bar{S}>0$ such that for $l$ large enough $S_{l,s}>\bar{S}$ and thus $e^{-lS_{l,s}}\leq e^{-l\bar{S}}.$ This, together with (79) ensures that ${\mathbb{E}}[P_{k,s}^{2}]$ is uniformly bounded and therefore, by Doob’s theorem, there exists $P_{\infty,s}$ such that $\lim_{k\rightarrow+\infty}P_{k,s}=P_{\infty,s}\enspace\text{almost surely}.$ Finally, from (76) and the convergence of $Q_{k,s}$ and $P_{k,s}$ we have, almost surely for every $s=1,\dots,d$,
@@
Since Th. 3.3 yields $\lim_{k\rightarrow+\infty}\|{x}^{i}_{k}-{x}^{j}_{k}\|=0$ a. s. for every $i,j=1,\dots,N$, we conclude that there exists ${x}_{\infty}\in\mathbb{R}^{d}$ such that ${x}^{i}_{\infty}={x}_{\infty}$ for every $i=1,\dots,N$, and we have the thesis. ∎
@@
$(\theta^{i}_{k}{x}^{i}_{k+1}+(1-\theta^{i}_{k}){x}^{i}_{k}-\hat{x}^{\alpha}_{k})=((1-\theta^{i}_{k}\gamma)I_{d}-\theta^{i}_{k}{\bf\Xi}_{k}^{i})\left(x^{i}_{k}-\hat{x}^{\alpha}_{k}\right)$ (82)
@@
By definition of $\Gamma_{A}$ and applying (iv) in Lemma 3.2, we get the thesis.
@@
for every $k=0,\dots,\bar{k}-1$ where $\bar{k}$ is such that (56) holds. This implies the thesis if $\bar{k}>k_{*}$ or if $\bar{k}\leq k_{*}$ and ${\mathbb{E}}[V_{k}]\leq\varepsilon$. Therefore the only thing left to show is that the remaining case, $\bar{k}\leq k_{*}$ and ${\mathbb{E}}[\|{x}^{\alpha}_{\bar{k}}-{x}^{*}\|^{2}]>C^{2}_{\bar{k}}$, cannot happen. Proceeding as in 63 and using Lemma 3.14 we have
@@
Proceeding as in the proof of Theorem 3.15, it is easy to see that if $\alpha=\hat{\alpha}$ sufficiently large and $M_{v}$ sufficiently small, the inequality above implies that ${\mathbb{E}}[\|\hat{x}^{\alpha}_{\bar{k}}-{x}^{*}\|^{2}]\leq C_{\bar{k}}^{2}.$

### v2

@@ 3.5 Convergence in expectation
$\mathbb{P}\left\{|{\cal I}_{\bar{k},r}|>0\right\}\geq 1-\delta\enspace\text{with}\ \delta\leq\hat{\delta}e^{-4\alpha M_{f}}$ (68)
where $\bar{k}$ is given in (56), $M_{f}$ is the upper bound from Corollary 3.7, and $\hat{\delta}\in(0,e^{4\alpha M_{f}})$. Then, for $\alpha$ sufficiently large and $M_{v}$ sufficiently small, it holds ${\mathbb{E}}[V_{k_{*}}]\leq\varepsilon$, with $k_{*}$ given in (57). Moreover, ${\mathbb{E}}[V_{k}]\leq\mu^{k}{\mathbb{E}}[V_{0}]$ for $k=0,\dots,k_{*},$ with $\mu$ given in (55).
@@
Remark 3.23.
From the theoretical analysis that we carried out, we can notice a trade-off between the exploration power of the system of particles and the convergence to a consensus point. Intuitively, if the number of particles $N$ is larger, they will visit more points in the search space, while it will be harder for the system to achieve consensus. This can be seen more rigorously in Theorems 3.5, 3.20 and 3.21 as condition 58 is more likely to hold if the number of particles $N$ is large while to ensure consensus one needs ${\bf\vartheta}=1-\gamma+8\xi(\log(\sqrt{2}N))^{1/2}<1$, which is more restrictive as $N$ increases. In the numerical result section we will show that the condition ${\bf\vartheta}<1$ appears in practice to be too restrictive.
Assumption (58), or its probabilistic relaxation (68), are necessary for our analysis, however in [11] it has been shown that for the continuous algorithm the analogous of $|{\cal I}_{k,r}|$, namely $\rho_{t}(\mathcal{B}_{r})$, is guaranteed to be strictly positive for $t\in[0,T]$ for any $T$.
@@ 4 Numerical Results
In this section we investigate on the sensitivity of our approach to noise in the objective function evaluation. We first consider the minimization of the well known Rastrigin function with Gaussian noise on the objective function and then we analyze the behaviour of CBO applied to finite sum minimization problems employing subsampling estimators of the objective function.
@@ 4.1 Rastrigin Function
We can observe that the absolute Gaussian noise affects the quality of the solution as expected, but we manage to obtain an average error of the order of $10^{-3}$ (against $10^{-5}$ obtained in the noise-free case), even when an high noise level is present. We also underline that the method is quite insensitive to this type of noise as increasing $\sigma_{0}$ we observe only a slight decrease of the overall obtained accuracy. Regarding the relative Gaussian noise, we note that the noise does not affect the accuracy of the computed approximation for values of the noise level smaller or equal than $0.25$, while the method is not able to handle higher values of the noise. In case of the mixed noise we obviously observe the effect of both the absolute and relative noise. Overall the observed behaviour is consistent with our theoretical analysis. Indeed, under Assumption A5, Theorems 3.20 and 3.21 shows that global convergence is obtained provided that the the upper bound $t_{0}+t_{1}M_{f}^{2}$ on the noise level is sufficiently small. In case of relative error, the bound on $\sigma_{1}$ is more severe than the bound on $\sigma_{0}$ at particles $x$ at which $f$ takes large value.
@@ 4.2 Subsampling
For both choices of the parameters we can observe that the final accuracy is not greatly affected by the subsampling, up to $\ell=0.25$ in the first case and $\ell=0.05$ in the second case. For $\gamma=0.1$ and $\xi=0.0056$ (Table 5) the number of iterations is approximately constant for all considered values of $\ell$, which implies that the computational cost decreases significantly as the sampling size decreases. For $\ell=0.5$ the overall computational cost is halved, but the average accuracy is less than $0.5\%$ smaller. For $\gamma=0.01$ and $\xi=0.1$ (Table 6) the number of iterations increases as the sampling factor decreases, but the overall computational cost decreases. The method that employs $\ell=0.1$ is approximately 2.5 times less expensive than the method that uses the full sample, while the average accuracy is only around $1\%$ smaller. Comparing the results for the two choices of the parameters (namely $\gamma=0.1,\ \xi=0.0056$ and $\gamma=0.01,\ \xi=0.1$) we notice that, despite the fact that the number of particles $N$ is larger in the first case, the particles converge in a smaller number of iterations, and that the overall accuracy is lower, compared to that reached by the method that employs smaller $\gamma$ and larger $\xi.$ This confirms that the condition ${\bf\vartheta}<1$ from the theoretical analysis may be too restrictive and the method can perform better in practice for choices of the parameters that increase the exploration of the search space.
@@ Acknowledgements
The research that led to the present paper was partially supported by INDAM-GNCS through Progetti di Ricerca 2024 and by PNRR - Missione 4 Istruzione e Ricerca - Componente C2 Investimento 1.1, Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN) funded by the European Commission under the NextGeneration EU programme, project “Advanced optimization METhods for automated central veIn Sign detection in multiple sclerosis from magneTic resonAnce imaging (AMETISTA)”, code: P2022J9SNP, MUR D.D. financing decree n. 1379 of 1st September 2023 (CUP E53D23017980001), project “Numerical Optimization with Adaptive Accuracy and Applications to Machine Learning”, code: 2022N3ZNAX MUR D.D. financing decree n. 973 of 30th June 2023 (CUP B53D23012670006).
@@
[14, Lemma 3.3 ] If $M\in\mathbb{R}^{N\times N}$ is such that $\sum_{j=1}^{N}M_{ij}=m$ for any index $i=1,\dots,N$, then for every vector ${y}\in\mathbb{R}^{N}$ we have $\operatorname{diam}(My)\leq(m-\operatorname{erg}(M))\operatorname{diam}(y).$
@@
[14, Lemma 4.1] Given any matrix $M\in\mathbb{R}^{N\times N}$, the ergodicity coefficient satisfies the following inequality: $\operatorname{erg}(M)\geq-2\|M\|_{\infty}$ for any $M\in\mathbb{R}^{N\times N}$
@@
$(M^{s}_{k})_{ij}=\begin{cases}1-\gamma-\eta^{i}_{k,s}+(\gamma+\eta^{i}_{k,s})\hat{v}^{\alpha}_{k,j}&\text{if }i=j\\ (\gamma+\eta^{i}_{k,s})\hat{v}^{\alpha}_{k,j}&\text{if }i\neq j.\end{cases}$ (73)
@@
Replacing (75) and (76) into (74) we get the thesis. ∎∎
@@
$Q_{k,s}=\gamma\sum_{l=0}^{k}(x^{i}_{l,s}-\hat{x}^{\alpha}_{l,s}),\enspace P_{k,s}=\sum_{l=0}^{k}(x^{i}_{l,s}-\hat{x}^{\alpha}_{l,s})\eta_{l,s}^{i}.$
@@
Since $\lim_{i\rightarrow\infty}S_{l,s}\geq 1-{\bf\vartheta}$ almost surely, there exist two stochastic variables $C_{1},C_{2}>0$ such that
@@
$={\mathbb{E}}\left[\sum_{l=0}^{k}(x^{i}_{l,s}-\hat{x}^{\alpha}_{l,s})^{2}(\eta_{l,s}^{i})^{2}+\sum_{l=0}^{k-1}\sum_{\begin{subarray}{c}p=0\\ p\neq l\end{subarray}}^{k-1}(x^{i}_{l,s}-\hat{x}^{\alpha}_{l,s})(x^{i}_{p,s}-\hat{x}^{\alpha}_{p,s})\eta_{l,s}^{i}\eta_{p,s}^{i}\right].$
@@
Since $\lim_{l\rightarrow+\infty}S_{l,s}\geq 1-{\bf\vartheta}>0$ almost surely, then there exists a constant $\bar{S}>0$ such that for $l$ large enough $S_{l,s}>\bar{S}$ and thus $e^{-lS_{l,s}}\leq e^{-l\bar{S}}.$ This, together with (80) ensures that ${\mathbb{E}}[P_{k,s}^{2}]$ is uniformly bounded and therefore, by Doob’s theorem, there exists $P_{\infty,s}$ such that $\lim_{k\rightarrow+\infty}P_{k,s}=P_{\infty,s}\enspace\text{almost surely}.$ Finally, from (77) and the convergence of $Q_{k,s}$ and $P_{k,s}$ we have, almost surely for every $s=1,\dots,d$,
@@
Since Th. 3.5 yields $\lim_{k\rightarrow+\infty}\|{x}^{i}_{k}-{x}^{j}_{k}\|=0$ a. s. for every $i,j=1,\dots,N$, we conclude that there exists ${x}_{\infty}\in\mathbb{R}^{d}$ such that ${x}^{i}_{\infty}={x}_{\infty}$ for every $i=1,\dots,N$, and the thesis follows. ∎∎
@@
$(\theta^{i}_{k}{x}^{i}_{k+1}+(1-\theta^{i}_{k}){x}^{i}_{k}-\hat{x}^{\alpha}_{k})=((1-\theta^{i}_{k}\gamma)I_{d}-\theta^{i}_{k}{\bf\Xi}_{k}^{i})\left(x^{i}_{k}-\hat{x}^{\alpha}_{k}\right).$ (83)
@@
By definition of $\Gamma_{A}$ and applying (iv) in Lemma 3.4, we get the thesis. ∎∎
Proof of Lemma 3.19.
$\|{x}^{\alpha}_{k}-{x}^{*}\|$ $=\left\|\frac{1}{\sum_{i=1}^{N}w_{i,k}^{\alpha}}\sum_{i=1}^{N}w^{\alpha}_{i,k}({x}_{k}^{i}-{x}^{*})\right\|\leq\frac{1}{\sum_{i=1}^{N}w_{i,k}^{\alpha}}\sum_{i=1}^{N}w^{\alpha}_{i,k}\left\|{x}_{k}^{i}-{x}^{*}\right\|$ (86)
By assumption A6 and the definition of $\tilde{r}$ we have
@@
for every $k=0,\dots,\bar{k}-1$ where $\bar{k}$ is such that (56) holds. This implies the thesis if $\bar{k}>k_{*}$ or if $\bar{k}\leq k_{*}$ and ${\mathbb{E}}[V_{k}]\leq\varepsilon$. Therefore the only thing left to show is that the remaining case, $\bar{k}\leq k_{*}$ and ${\mathbb{E}}[\|{x}^{\alpha}_{\bar{k}}-{x}^{*}\|^{2}]>C^{2}_{\bar{k}}$, cannot happen. Proceeding as in (63) and using Lemma 3.19 we have
@@
Proceeding as in the proof of Theorem 3.20, it is easy to see that if $\alpha=\hat{\alpha}$ sufficiently large and $M_{v}$ sufficiently small, the inequality above implies that ${\mathbb{E}}[\|\hat{x}^{\alpha}_{\bar{k}}-{x}^{*}\|^{2}]\leq C_{\bar{k}}^{2}.$ ∎∎